# SPECIAL RELATIVITY

### v9

@@ 8.1 Potential of Earth-Based Electron Microscopes
Above we discussed the new relativistic particle kinematics for rotating frame of references. Since we live on rotating earth frame, difference in relativistic kinematics between rotating and non-rotating frames of references is of practical as well as theoretical significance. We propose an aberration of particle experiment using an earth-based particle source. Experiments can be explained on the basis of special relativity. A correct solution of this problem in the earth-based frame requires the use of metric tensor even in first-order experiments since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a relativistic particle beam propagation.
@@
We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source. One could use a vertically oriented optical setup. We suppose that the electron beam is imaged by a lens to a spot which lies in the image plane. Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement. Due to the high stability, the aberration increment could be observed (1).
@@
Consider the inertial sun-based reference system. In this system there is the earth which rotates around the sun with orbital velocity $v$. Developing into powers of $v/c$ we can classify effects for velocities $v\ll c$ as of the first order, of the second order or of higher orders. The essentially relativistic effects are of the second (or higher) orders. Clearly, in the case of the rotating (around the sun) earth-based frame, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$.
@@
Suppose that an earth-based observer performs an aberration of particle measurement. The aberration angle varies with one-year temporal period. An approximate formula to express the aberration angle can be found by using Langevin metric in the earth-based frame or by using Wigner rotation theory in the sun-based frame. In the aberration of particles, the aberration angle is the apparent angular deviation of the position of the particle source relative to the real location of the source. The reference axis in the earth-based frame can be formed by plumb line, which is the most fundamental local earth-based coordinate system.
@@
When measured with an earth-based particle source, the aberration angle is $\vec{\theta}_{a}=-\vec{\theta}_{w}=-(1-1/\gamma_{p})\vec{V}\times\vec{v}/V^{2}$, where $v$ is the orbital velocity of the earth relative to the sun, $V$ is the particle velocity, and $\gamma_{p}=1/\sqrt{1-V^{2}/c^{2}}$ is the particle relativistic factor. This means that the angle of Wigner rotation depends on $V^{2}/c^{2}$. It should be note that presented formula can be considered only as a first order approximation in $v/c$. In the theory of (infinitesimal) Wigner rotation we must also consider the relativistic parameter $V/c$. If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$.
@@
Let us calculate the parameters for the STEM setup. The Wigner rotation angle parameter (constant of aberration) defines the scale of the aberration increment, and is equal to $\theta_{w}=(1-1/1/\gamma_{p})v/V\simeq 40\mu\mathrm{rad}$. Here $\gamma_{p}\simeq 1.39$, $V/c=0.7$, $v/c\simeq 100\mu\mathrm{rad}$. Earth moves around the sun, with a consequent change of the direction of $\vec{v}$. Therefore the angle of aberration changes. The total change of the order of $40\mu\mathrm{rad}$. It is oscillatory with a period of one year.
@@
A specific aspect of our case study needs further investigations. The proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. Obviously, it is important that observation should be recorded in the shortest possible time. In principle, records could be taken over a period of one day. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth rotation axis (2). In practice, the measurement procedure is complicated by a number of factors, Fig. 42. A corrections has to be made for the observer position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$. The image appears to move in an ellipse. The shape of the aberration ellipse obviously depends on the observer’s latitude $\phi$. For an observer on the equator the ellipse degenerates into a line segment and for an observer at the pole of the earth, the ellipse is a circle. At the value $l=2$ mm of the focal length (3), the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm. The value of minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm. When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and various from a value about 10 nm/60 min to about 20 nm/60 min through the day (4).
A large number of incorrect expression for the Wigner rotation (it is often called the ”Thomas rotation”) can be found in the literature. For those who have already received knowledge about the Wigner rotation from the well-known textbooks, it may be expedient to begin its aberration of particle (theoretical and experimental) development with the microscopic approach. The point is that one can compute any relativistic quantity directly from the underlying theories of matter without involving relativity at all. Wigner rotation associated with the transformation from the inertial (sun-based) coordinate system to the rotation (earth-base) coordinate system may be regarded as a result of the action of certain force. This problem has no relation to the problem of the ether and is already formulated within the framework of special relativity. The electromagnetic forces which govern the properties of an emitted electron beam must be affected by (earth-based) source acceleration around the sun in a such a way that they lead to a deviation of the electron transport direction. In fact, there is a machinery behind the electron momentum $\vec{p}$ changes when the source and observer (with his measuring instruments) are at rest in the rotating frame. Its origin is explained in the framework of the electrodynamics theory.
We can obtain the electrodynamics equations in rotating reference frame using the Galilean transformation. When we consider the electrodynamics equations under Galilean transformation, there is a question: what is the transformation law for the electromagnetic fields $\vec{E}$ and $\vec{B}$? The fields $\vec{E}_{n}$ and $\vec{B}_{n}$ observed in the accelerated frame where the particle source is at rest are not the same as those $\vec{E}$ and $\vec{B}$ in the inertial frame before the active Galilean boost.
We need to find the relation between $\vec{E}_{n}$ and $\vec{B}_{n}$ and $\vec{E}$ and $\vec{B}$. To do this, we transform coordinates $(t,x)$ that would be coordinates of an inertial observer $S$ moving with velocity $-v$ with respect to the observer $S_{n}$, using a Galilean transformation: we substitute $x_{n}=x-vt$, while leaving time unchanged $t_{n}=t$ into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain Langevin metric $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. We can obtain the electrodynamics equations in rotating frame using the Galilean transformation (with velocity $-v$) of the Maxwell’s equations. With the Galilean transformation, the field transformation, for the case $\vec{B}=0$, is $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. This expression describe first order (in $v/c$) effects only. Here $\vec{v}$ is the velocity vector of the earth in the sun-based frame.

### v10

@@ 10.1 Potential of Earth-Based Electron Microscopes
Above we discussed the new relativistic particle kinematics for rotating frames of reference. Since we live on a rotating earth frame, the difference in relativistic kinematics between rotating and non-rotating frames of reference is of practical as well as theoretical significance. We propose an aberration of particle experiment using an earth-based particle source. Experiments can be explained based on special relativity. A correct solution to this problem in the earth-based frame requires the use of metric tensor even in first-order experiments.
@@
We propose an experiment using commercial 200-kV scanning transmission electron microscope (STEM) as a particle source. One could use a vertically oriented optical setup. We suppose that the electron beam is imaged by a lens to a spot that lies in the image plane. Measurement of the spot shift with respect to the optical axis (which is parallel to the zenith-nadir axis) is equivalent to the determination of the angular displacement. Due to the high stability, the aberration increment could be observed  [^42] The dedicated STEM with high stability was made [47]. It was developed based on a 200-kV commercial instrument. The electron source was installed in the anti-seismic room. Temperature fluctuations in the room were within 0.2 ${}^{o}C$. The magnetic field in the room was less than $0.2mG$ near the source column. To prevent thermal drift, a shield box was placed over the high-voltage generator. The column was covered with rubber sheets to reduce the temperature fluctuations. The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$. The drift speed and direction during this period were approximately constant. The average drift speed was 0.12 nm/min. It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min..
@@
Consider the inertial sun-based reference system. In this system, there is the earth which rotates around the sun with orbital velocity $v$. Developing into powers of $v/c$ we can classify effects for velocities $v\ll c$ as of the first order, of the second order, or higher orders. Clearly, in the case of the rotating (around the sun) earth-based frame, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$.
@@
Suppose that an earth-based observer performs an aberration of particle measurement. The aberration angle varies with a one-year temporal period. An approximate formula to express the aberration angle can be found by using the Langevin metric in the earth-based frame or by using the Wigner rotation theory in the sun-based frame. In the aberration of particles, the aberration angle is the apparent angular deviation of the position of the particle source relative to the real location of the source. The reference axis in the earth-based frame can be formed by a plumb line, which is the most fundamental local earth-based coordinate system.
@@
When measured with an earth-based particle source, the aberration angle is $\vec{\theta}_{a}=-\vec{\theta}_{w}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{v}$ is the orbital velocity of the earth relative to the sun, $\vec{V}$ is the velocity of the accelerated electron, and $\gamma=1/\sqrt{1-V^{2}/c^{2}}$ is the particle relativistic factor. This means that the angle of Wigner rotation depends on $V^{2}/c^{2}$. It should be noted that the presented formula can be considered only as a first-order approximation in $v/c$. In the theory of (infinitesimal) Wigner rotation, we must also consider the relativistic parameter $V/c$. If a particle motion velocity is non-relativistic, the binomial expansion yields $\vec{\theta}_{a}=-[V^{2}/(2c^{2})]\vec{V}\times\vec{v}/V^{2}$.
@@
Let us calculate the parameters for the STEM setup. The Wigner rotation angle parameter (constant of aberration) defines the scale of the aberration increment and is equal to $\theta_{w}=(1-1/1/\gamma)v/V\simeq 40$ $\mu\mathrm{rad}$. Here $\gamma\simeq 1.39$, $V/c=0.7$, $v/c\simeq 100$ $\mu\mathrm{rad}$. Earth moves around the sun, with a consequent change of the direction of $\vec{v}$. Therefore the angle of aberration changes. The total change of the order of $40\mu\mathrm{rad}$. It is oscillatory with a period of one year.
@@
A specific aspect of our case study needs further investigation. The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. Obviously, it is important that observation should be recorded in the shortest possible time. In principle, records could be taken over a period of one day. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis 434343 The constant of diurnal aberration defines the scale of the aberration increment and is equal to $\theta_{d}=(1-1/1/\gamma)v_{e}/V\simeq 0.6$ $\mu\mathrm{rad}$. Here $\gamma\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55$ $\mu\mathrm{rad}$ where $v_{e}=0.45$ km/s is the linear velocity of the earth rotation at equator.. In practice, the measurement procedure is complicated by a number of factors, Fig. 40. A correction has to be made for the observer’s position on the earth’s surface and the obliquity of the ecliptic $\delta=23^{o}$. The image appears to move in an ellipse. The shape of the aberration ellipse depends on the observer’s latitude $\phi$. For an observer on the equator, the ellipse degenerates into a line segment, and for an observer at the pole of the earth, the ellipse is a circle. At the value $l=2$ mm of the focal length  [^44] Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column. The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen. Electrons entering the lens column appear to come from a virtual source. The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm. Using a virtual electron source as its object, C1 produces a real image. Because the lens is located $D_{1}=15$ cm below the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$. The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to vary continuously over a wide range. The distance between lens centers is typically 10 cm. Spacemen is located $D_{3}=25$ cm below the C2., the major axis of the aberration ellipse of the order of 80 ($\cos\delta)$ nm. The value of the minor axis depends on the observer’s position and is given by 80 ($\cos\delta\sin\phi))$ nm. When the latitude is $\phi=35^{o}$, the value of the scale of image shift per hour depends on the local time and varies from a value of about 10 nm/60 min to about 20 nm/60 min through the day  [^45] Experimental results (see Fig. 44 p.167 in [47]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$. It is important to note that we used the typical focal length (2 mm) in our estimations. The details of the optical setup may influence the image shift. .
We want now to discuss a simple scaling model for the stellar aberration, Fig. 41. Let us consider the scale transformation of the source linear dimension and the distance between the source and observer. We will obtain a condition for optical similarity between the aberration of light from a distant star and from the earth-based source. This scale transformation may help in the initial design of the earth-based setup using the incandescent lamp as a source instead of a star.
Figure 41: Scaling model for the stellar aberration. Earth-based setup using incandescent lamp. We know (without using our knowledge of special relativity) what would happen from the astronomical observations. Without looking at anything external to the earth, one could determine the speed of the earth around the sun using aberration of light measurements.
A star ( or an incandescent lamp) is a system of elementary statistically independent point sources. Point-source field diffraction can be divided into categories - the Fresnel (near-zone) diffraction and Fraunhofer ( or telescopic) diffraction. If the Fraunhofer (or equivalently, far zone) approximation $z>D^{2}/\lambda$ is satisfied, then the quadratic phase factor is approximately unity over the detection aperture. Here $D$ is the aperture width, $\lambda$ is the wavelength of the radiation, and $z$ is the distance between the source and observer. In the far zone, a point source produces in front of pupil detection effectively a plane wave. The plane wavefronts of starlight entering the telescope are imaged by the lens to a diffraction spot that lies in the focal plane. It is worth noting that we consider the divergence of the transmitted radiation, $\lambda/D$, that is relatively small compared to the aberration angle, $\theta_{a}=v/c$.
An incandescent lamp is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. In our case of interest, imaging system is situated in the far zone $D^{2}/(\lambda z)<1$. It is the source linear dimension $d$ (source diameter) that determines the coherent area of the observed radiation $(z\lambda/d)^{2}$. The condition for neglecting the transverse size of the source, $d$, formulated as the requirement that the change of the correlation function along the aperture be less than unity, that is $d<D$ 464646 In all cases of practical interest, telescopes are situated in the far zone of the distant stars. If the far zone approximation $z>d$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form. For example, the coherence area of the light emitted by the circular incoherent source is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$. The minimal coherent area of light observed on the earth’s surface from the nearest star has a diameter of about $\lambda z/d\simeq$ 10 m. This diameter is larger than a typical telescope aperture $D$. In this situation, a star can be considered as a point source and a star image is a point spread function in the image plane of a telescope. The star consists of an extended collection of independent ”point” sources. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$ and far zone condition $z>D^{2}/\lambda$ for the elementary ”point” source is automatically satisfied. Our scaling model is based on the far zone approximation. We have only to change the distance and the source diameter. With the limitation of the small parameters assumption ($Dd/(z\lambda)<1$, $D^{2}/(z\lambda)<1$) the proposed scaling method gives $d<D$..
Let us estimate the parameters for the earth-based setup using the incandescent lamp. We want to describe some thought - but possible - experiments. At optical frequencies, the conditions required for the validity of the Fraunhofer approximation can be severe ones. For example, at a wavelength of $\lambda=0.3$ $\mu$m and an aperture width of $D=3$ cm, the observation distance must satisfy $z>3000$ m. Obviously, the required conditions are met in a number of important problems. No one has ever done such an experiment, but we know (without using our knowledge of special relativity) what would happen from the astronomical observations. The purpose of this thought experiment is to suggest that without looking at anything external to the earth, one could determine the speed of the earth around the sun using aberration of light measurements.
It should be noted that Fraunhofer diffraction patterns can be observed at distances much closer than implied by equation $z>D^{2}/\lambda$ provided the aperture is illuminated by a spherical wave converging towards the observer, or if a positive lens is properly situated between the observer and the aperture.
A working way of the earth-based optical setup to detect the aberration of light is schematically shown in Fig. 42. This is the well-known two-lens image formation scheme. This scheme allows for magnification by changing the focal distance of the second lens but for simplicity, in the following, we will assume that the two focal distances are the same (i.e. we consider 1:1 imaging). This two-lens setup is usually employed for image-processing purposes, as it can be better used for image modification than the single-lens system. Given the two-lens setup discussed above, we consider the relatively simple problem of characterization of point source radiation in the image plane. We assume the two lenses in Fig. 42 are identical.
To perform an aberration of light measurement, the source linear dimension $d$ should be radically decreased. We assume that the order of magnitude of the dimensions of the ”point” source is about $\lambda$, where $\lambda$ is the optical radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. In the optics for such ”point” source emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as an ensemble of atoms) by quantum mechanics. With the help of modern lithography, it is not difficult to produce an unresolved point source at optical wavelengths and let sufficiently bright radiation from it  [^47] It may be possible to use the masked spatially incoherent UV light source. One could use the light source onto a (micron diameter) hole of a specially designed mask. The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source..
Consider a diffracting aperture that is circular, and let the radius of the aperture be $a$. The physical properties of the lens can be combined in a single number $f$ called the focal length. We assume that the image is produced by a diffraction-limited optical system (i.e. a system that is free from aberrations). Once the wavelength is fixed, the resolution only depends on the numerical aperture $NA=a/f\ll 1$ of the system. The response of the system at point $(x_{i},y_{i})$ of the image plane to a $\delta$ function input at coordinates $(x_{o},y_{o})$ of the object plane is called the point spread function of the system. In our case of interest, one obtains the following amplitude point spread function: $A(x_{i},y_{i})=J_{1}(\alpha)/\alpha$, which is the diffraction pattern of a circular aperture, where $\alpha=2\pi ar_{i}/(f\lambda)$, $x_{0}=0,y_{0}=0$, $r_{i}=(x_{i}^{2}+y_{i}^{2})^{1/2}$. The first zero of the Airy’s pattern is at $r_{i}=0.6\lambda/NA$.
Let us estimate the parameters for the optical setup shown in Fig.42. Suppose $a=1.5$ cm and $\lambda=0.3$ $\mu$m. We will use $NA=0.1$. The resolution analysis is reduced to the theoretical framework of standard imaging theory. One can take advantage of well-known resolution criteria like Rayleigh criteria i.e. $20.6\lambda/NA=$ 2 $\mu$m. One could use a vertically oriented optical setup. The constant of aberration defines the scale of the aberration increment, and is equal to $v/c=100$ $\mu$rad. Measurement of the spot shift with respect to the optical axis (which is parallel to the nadir direction) is equivalent to the determination of the angular displacement. As already pointed out, the proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. In principle, records could be taken over a period of one day. The image appears to move in an ellipse. The major axis of the ellipse is about 14 $\mu$m. The value of the minor axis depends on the observer’s position and is given by $14(\sin\phi)$ $\mu$m  [^48] The important question is that of light detection. The small image size form rather sever restriction on the use of the CCD as a detector. In fact, for the case considered numerically image size is near 2 $\mu$m. One possibility to overcome this practical difficulty is photon detection with an MCP (microchannel plate photomultiplier) detector. It may be possible to use a (commercially available) precision 2D translation stage with a resolution of 0.5 $\mu$m to control the motion of a hole in the opaque screen. The aperture diameter in this example may be set to 1 micron or smaller. The screen is fixed horizontally on the translation stage which is placed in the image plane. The MCP-PMT may realize the measurement of (transmitted through a hole) photons. The proposed design has the advantage of device compactness. The height of our proposed optical configuration is about one meter..
We derived our results under the assumption that the order of magnitude of the dimensions of the source is about $\lambda$. It should be noted that the resolution of proposed optical techniques is not limited by the size of the incoherent source. The radiation field generated by an incoherent source can be seen as a linear superposition of fields of individual elementary point sources. The image of an elementary point source is a point spread function. An optical system using incoherent light could be completely described in therm of convolution integral relating the object distribution to the image distribution. In such a description, the optical system is characterized by its impulse response, i.e. the point image. The rounding or softening of the source’s edges is characteristic of the convolution (”smoothing”) integral.
It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like oscillatory image shift within a period of one day. This argument gives the reason why our technique can be extended to allow for the reconstruction of the angle of aberration from a measurement of image shift of an incoherent source. The edge object has become useful in aberration shift detection. An alternative application of a ”point” source is an incoherent source with (physically) sharp edges. The edge width of the source is strictly related to the practical resolution achievable. For our particular example, the edge width should be about $1\mu$m.
11 The Aberration of Light from a Laser Source
11.1 Theory of a Laser with Optical Resonator
Suppose that an earth-based observer performs an aberration of light measurement using a laser as a light source. Let us stress that the earth’s orbital motion cannot be measured using a laser. To understand this phenomenon, we consider the simple case when a laser resonator is equipped with plane mirrors. For oscillation to occur, the total loss in power due to diffraction and reflection loss must be less than the power gained by travel through the active medium. Thus diffraction loss is expected to be an important factor, both in determining the start-oscillation condition and in determining the distribution of energy in the interferometer during oscillation. The purpose of our study is to investigate the effect of diffraction on the electromagnetic field in a Fabry-Perot interferometer in free space. The conclusion can be applied equally well to gaseous lasers provided the interferometer is immersed in the active medium.
The electric field $\vec{E}$ of an electromagnetic wave in a passive resonator satisfies the equation $\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$. For simplicity, we shall assume that the field in the resonator is circularly polarized. Such an assumption does not violate the generality of the analysis because the polarization degeneracy takes place in the system under study. The electric field vector in the plane resonator is presented as an oscillation superposition with different longitudinal wave numbers, $E_{x}+iE_{y}=\sum_{m}\widetilde{E}_{m}(x,y,t)\exp(-i\omega_{m}t)\sin(K_{m}z+\delta)$, where $K_{m}=m\pi/l-i/(nl)$, $\omega_{m}=m\pi c/l$, $\delta=i/n$, $n$ is the (complex) refractive index of the mirror matter ($|n|\gg 1)$, $l$ is the distance between the mirrors, and $m$ is an integer number ($m\gg 1$). It is evident that this expression satisfies Leontovich’s boundary conduction  [^49] We have presented an example of introducing approximate boundary conditions. These approximate boundary conditions were derived by Leontovich. They are valid on the surface of a material having a large value of the modulus of the refractive index. An extended theoretical treatment of Leontovich’s boundary conditions appears in the monograph [48]. on the mirror surface when $z=0$ and $z=l$. We assume the field change per one resonator pass to be small i.e. $|\partial\widetilde{E}_{m}/\partial t|l/c\ll|\widetilde{E}_{m}|$, which means that the laser frequency is close to the natural frequencies $\omega_{m}$ of the longitudinal resonator modes. Let us find a solution for the field amplitude $\widetilde{E}_{m}$ in the form $\widetilde{E}_{m}=\Phi(x,y)\exp(\Lambda t)$. Substituting the ansatz in the wave equation we get $\nabla_{\perp}\Phi+2i\omega_{m}[\Lambda/c^{2}+2/(nl)]\Phi=0$.